# Tensor Product Attention Is All You Need

### v3

@@ A.8 Decoding Speed during Inference Time of MHA, MQA, GQA, MLA, and TPA
[…]
Thus, in this setup, TPA can significantly reduce the flops needed for computing the $\mathbf{Q}(\mathbf{x}_{T})\bm{K}^{\top}$ operation at each decoding step. The actual end-to-end wall-clock speedup also depends on kernel fusion, caching strategies, and hardware implementation details, but the factorized formulation offers a pathway to more efficient decoding than standard attention.
Appendix B Higher-Order Tensor Product Attention
All prior discussions have focused on a second-order factorization in which each rank-$R_{Q}$ (and similarly $R_{K}$, $R_{V}$) component is the outer product of two vectors: one in $\mathbb{R}^{h}$ (the “head” dimension) and one in $\mathbb{R}^{d_{h}}$. We now generalize this by introducing an additional latent factor, yielding a third-order (or higher) factorization reminiscent of canonical polyadic (CP) decomposition. Concretely, for a single token $t$, we write
@@ Appendix B Higher-Order Tensor Product Attention
where the newly introduced factor $\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{\mathbf{c}}}$ can be viewed as a learnable gate or modulation term. Analogous expansions apply to $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$. In practice, these triple (or higher-order) products still collapse into a matrix in $\mathbb{R}^{h\times d_{h}}$. One straightforward way to achieve this collapse is to split the feature dimension $d_{h}$ such that $d_{b}\times d_{c}=d_{h}$,
$\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}},\quad\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}},\quad\operatorname{vec}\bigl{(}\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\bigr{)}\in\mathbb{R}^{d_{h}}.$
This additional factor can enhance expressiveness without necessarily increasing the base rank. Conceptually, it can act as a learnable nonlinearity or gating mechanism. One could also tie or share $\mathbf{c}^{Q}_{r}$ across queries, keys, and values, to reduce parameter overhead.
A similar setup holds for keys (with rank $R_{K}$) and values (with rank $R_{V}$). Although this extra dimension adds to the parameter count, it can reduce the required rank to achieve a certain level of representational power.
From a memory perspective, higher-order TPA still leverages factorized KV caching: only the factors $\mathbf{a}(\mathbf{x}_{t}),\mathbf{b}(\mathbf{x}_{t})$, and $\mathbf{c}(\mathbf{x}_{t})$ for each past token are cached. As usual, a trade-off arises between model capacity and the overhead of memory and computing. Nonetheless, moving from a rank-$\bigl{(}R_{Q},R_{K},R_{V}\bigr{)}$ matrix factorization to a higher-order tensor decomposition can provide additional flexibility and increased capacity.
B.1 RoPE Compatibility in Higher-Order TPA
Rotary positional embeddings (RoPE) remain compatible even under higher-order factorizations. In second-order TPA, RoPE can be treated as an invertible blockwise linear map acting on the last dimension of $\mathbf{Q}_{t}$ or $\mathbf{K}_{t}$. The same argument carries over when a third factor $\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})$ is present. Suppose RoPE acts on the $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})$ portion (of dimension size $d_{b}$), we have the following theorem.
Theorem 2 (RoPE Compatibility in Higher-Order TPA).

### v4

@@ C.8 Inference Time Decoding of Different Attention Mechanisms
TPA Total: $\quad 1568M+2080M=3648M$
For MQA and GQA, assuming $H=32$ query heads and $d_{h}=64$:
MQA/GQA Total: $\quad 2MHd_{h}=2\cdot M\cdot 32\cdot 64=4096M$
For MLA, with $H=32,d_{c}=256,d_{h}^{R}=32$:
MLA Logits: $\quad MH(d_{c}+d_{h}^{R})=M\cdot 32\cdot(256+32)=M\cdot 32\cdot 288=9216M$
MLA Value Agg.: $\quad MHd_{c}=M\cdot 32\cdot 256=8192M$
MLA Total: $\quad M\cdot 32\cdot(2\cdot 256+32)=M\cdot 32\cdot 544=17408M$
In this configuration, TPA requires fewer FLOPs ($3648M$) than MHA ($4096M$). The FLOPs for scaled logits in TPA ($1568M$) are less than MHA’s ($2048M$). The value aggregation FLOPs for TPA ($2080M$) are comparable to MHA’s ($2048M$).
This reduction in FLOPs, particularly when $R_{K}$ and $R_{V}$ are small (e.g., 1), combined with optimized memory access patterns in algorithms like FlashTPA Decoding, contributes to TPA’s competitive decoding speed, especially for long sequences. The actual end-to-end wall-clock speedup also depends on kernel fusion, hardware specifics, and the efficiency of einsum implementations, but the factorized formulation offers a clear path to reduced computational load.
Appendix D More on FlashTPA Decoding Algorithm
This section provides further details on the FlashTPA Decoding algorithm (see Algorithm 1), including the setup of factorized components.
Factorized Component Setup for FlashTPA Decoding. Let $B$ be the batch size, $N$ the number of query tokens (which is 1 for decoding, $N=1$), $M$ the current length of the KV cache, $H$ the number of attention heads, and $R_{Q}$ the rank of the query factorization. The feature dimensions for the factorized components are $D$ (for query $\mathbf{B}_{Q}$ and key $\bm{b}^{K}$) and $E$ (for value $\bm{b}^{V}$). Typically, $D$ and $E$ correspond to the head dimension $d_{h}$.
The query for the current token $\mathbf{x}_{t}$ (where $t$ is the current time step, $N=1$) is factorized as $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$, where $\mathbf{A}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{H\times R_{Q}}$ (derived from the input tensor ‘Aq’ of shape $(B,1,H,R_{Q})$ for the current token) and $\mathbf{B}_{Q}(\mathbf{x}_{t})\in\mathbb{R}^{R_{Q}\times D}$ (similarly from ‘Bq’ of shape $(B,1,R_{Q},D)$).
The cached keys $\mathbf{K}_{s}$ and values $\mathbf{V}_{s}$ for past tokens $s\in[1,M]$ are stored in their factorized form. For clarity in the data flow diagram (Figure 2 in the main paper) and the pseudo-code in Algorithm 1, we present the case where cached keys and values use ranks $R_{K}=1$ and $R_{V}=1$. The general TPA formulation allows for $R_{K}>1$ and $R_{V}>1$ for cached items (by storing, for example, $\mathbf{A}_{K}(\mathbf{x}_{s})\in\mathbb{R}^{R_{K}\times H}$ and $\mathbf{B}_{K}(\mathbf{x}_{s})\in\mathbb{R}^{R_{K}\times D}$). The einsum operations detailed in Algorithm 1 can be extended to accommodate higher ranks for keys and values by including summations over these additional rank dimensions. For $R_{K}=1$ and $R_{V}=1$, the cached $s$-th key and value are: $\mathbf{K}_{s}=\mathbf{a}^{K}(\mathbf{x}_{s})\otimes\bm{b}^{K}(\mathbf{x}_{s})\quad\in\mathbb{R}^{H\times D}$, $\mathbf{V}_{s}=\mathbf{a}^{V}(\mathbf{x}_{s})\otimes\bm{b}^{V}(\mathbf{x}_{s})\quad\in\mathbb{R}^{H\times E}$. Here, $\mathbf{a}^{K}(\mathbf{x}_{s})\in\mathbb{R}^{H}$ (from the cached tensor $\mathbf{a}^{K}_{\text{cache}}$ of shape $(B,M,H)$), $\bm{b}^{K}(\mathbf{x}_{s})\in\mathbb{R}^{D}$ (from $\bm{b}^{K}_{\text{cache}}$ of shape $(B,M,D)$), $\mathbf{a}^{V}(\mathbf{x}_{s})\in\mathbb{R}^{H}$ (from $\mathbf{a}^{V}_{\text{cache}}$ of shape $(B,M,H)$), and $\bm{b}^{V}(\mathbf{x}_{s})\in\mathbb{R}^{E}$ (from $\bm{b}^{V}_{\text{cache}}$ of shape $(B,M,E)$).
D.1 Detailed Computation Steps of FlashTPA Decoding Algorithm
The algorithm proceeds through a series of einsum operations, which correspond to specific tensor contractions:
  1. 1.
    Compute $\mathbf{B}_{Q}$-$\bm{b}^{K}$ inner products ($S^{(1)}$): For each query (batch $b$, query index $n=0$), each query rank component $r$, and each cached key position $m$, we compute the dot product of the $D$-dimensional feature vector of the $r$-th query factor $(\mathbf{B}_{Q})_{b,n,r,:}$ with the $D$-dimensional feature vector of the $m$-th cached key $(\bm{b}^{K}_{\text{cache}})_{b,m,:}$.
    $S^{(1)}_{b,n,m,r}=\sum_{d=1}^{D}(\mathbf{B}_{Q})_{b,n,r,d}\cdot(\bm{b}^{K}_{\text{cache}})_{b,m,d}.$
    This step captures the similarity between the feature parts of the query factors and the cached key factors.
  2. 2.
    Combine with query factor $\mathbf{A}_{Q}$ ($S^{(2)}$): The result $S^{(1)}$ is then combined with the head-specific query factors $\mathbf{A}_{Q}$. For each head $h$, the contributions from all query ranks $r$ are summed up.
    $S^{(2)}_{b,n,m,h}=\sum_{r=1}^{R_{Q}}(\mathbf{A}_{Q})_{b,n,h,r}\cdot S^{(1)}_{b,n,m,r}.$
    This yields an intermediate score for each query, cached key, and head.
  3. 3.
    Incorporate cached key factor $\mathbf{a}^{K}_{\text{cache}}$ for full logits ($\mathcal{L}$): The head-specific factor of the cached keys, $\mathbf{a}^{K}_{\text{cache}}$, is multiplied with $S^{(2)}$ to produce the final unscaled attention logits.
    $\mathcal{L}_{b,h,n,m}=S^{(2)}_{b,n,m,h}\cdot(\mathbf{a}^{K}_{\text{cache}})_{b,m,h}.$
    This logit $\mathcal{L}_{b,h,n,m}$ represents $s_{Q}^{-1}\cdot(\mathbf{Q}_{t}\mathbf{K}_{m}^{\top})_{h}$, i.e., the dot product between the query $\mathbf{Q}_{t}$ and the $m$-th cached key $\mathbf{K}_{m}$ for head $h$, scaled by $R_{Q}$ (since $s_{Q}=1/R_{Q}$). The ‘einsum‘ operation also rearranges dimensions to $(B,H,N,M)$ for the subsequent softmax operation.
  4. 4.
    Apply Scaling and Softmax ($\bm{\alpha}$): The logits $\mathcal{L}$ are scaled by the product of $s_{Q}$ (e.g., $1/R_{Q}$), $s_{K}$ (e.g., $1/R_{K}$, which is 1 if $R_{K}=1$), and $s_{\text{total}}$ (e.g., $1/\sqrt{D}$, where $D$ is the dimension used for dot products, typically $d_{h}$). Softmax is then applied across the $M$ cached key positions for each head independently.
    $\alpha_{b,h,n,m}=\operatorname{Softmax}_{m}\left(\mathcal{L}_{b,h,n,m}\cdot s_{Q}\cdot s_{K}\cdot s_{total}\right).$
    These are the attention probabilities.
  5. 5.
    Compute weighted sum with cached value factor $\mathbf{a}^{V}_{\text{cache}}$ ($\mathbf{O}^{(A)}$): The attention probabilities $\bm{\alpha}$ are used to weight the head-specific factors of the cached values, $\mathbf{a}^{V}_{\text{cache}}$.
    $\mathbf{O}^{(A)}_{b,n,m,h}=\alpha_{b,h,n,m}\cdot(\mathbf{a}^{V}_{\text{cache}})_{b,m,h}.$
    This step prepares the value components for the final aggregation. The ‘einsum‘ operation also reorders dimensions to $(B,N,M,H)$.
  6. 6.
    Incorporate cached value factor $\bm{b}^{V}_{\text{cache}}$ and apply final scaling ($\mathbf{O}$): Finally, the intermediate weighted value factors $\mathbf{O}^{(A)}$ are combined with the feature-specific factors of the cached values, $\bm{b}^{V}_{\text{cache}}$, by summing over the $M$ cached positions. The result is then scaled by $s_{V}$ (e.g., $1/R_{V}$, which is 1 if $R_{V}=1$).
    $\mathbf{O}_{b,n,h,e}=\left(\sum_{m=1}^{M}\mathbf{O}^{(A)}_{b,n,m,h}\cdot(\bm{b}^{V}_{\text{cache}})_{b,m,e}\right)\cdot s_{V}.$
    This produces the final output tensor of shape $(B,1,H,E)$.
D.2 Triton FlashTPA Decoding Kernel
We implement the experiments using Triton language [56], and the detailed implementation pseudo code is displayed in Algorithm 3. The Triton FlashTPA decoding kernel supports parallelism across the number of heads, rank, and head dimensions. And in 3, we will show the experiment results with this kernel.
Algorithm 3 Triton FlashTPA Decoding Kernel
1:Input Tensors: $\mathbf{A}_{Q}(B,N,H,R_{Q})$, $\mathbf{a}^{K}(B,M,H)$, $\mathbf{a}^{V}(B,M,H)$, $\mathbf{B}_{Q}(B,N,R_{Q},D)$, $\bm{b}^{K}(B,M,D)$, $\bm{b}^{V}(B,M,E)$
2:Scaling factors: $s_{\text{total}},s_{Q},s_{K},s_{V}$; Dimensions: $B,N(=1),M,H,R_{Q},D,E$
3:Kernel Block dims: $B_{H},B_{R},B_{D},B_{E}$; Sequence Blocking: $M_{\text{block}},M_{\text{chunk}}$
4:Program IDs: $p_{id_{B}},p_{id_{H}},p_{id_{M}}$
5:Partial Output $\mathbf{O}_{\text{partial}}(B,\text{Num}_{M},N,H,E)$, Log-Sum-Exp $\mathbf{LSE}_{\text{partial}}(B,\text{Num}_{M},H)$
6:
7:$b\leftarrow p_{id_{B}}$; $h_{\text{start}}\leftarrow p_{id_{H}}\cdot B_{H}$
8:$m_{\text{block}\_\text{start}}\leftarrow p_{id_{M}}\cdot M_{\text{block}}$; $m_{\text{block}\_\text{end}}\leftarrow\min((p_{id_{M}}+1)\cdot M_{\text{block}},M)$
9$\triangleright$ $B_{H},B_{R},B_{D},B_{E}$ are tile sizes for dimensions H, R, D, E respectively.
10:
11$\triangleright$ Initialize accumulators for the head block
12:$\mathbf{o}_{\text{accum}}\leftarrow\mathbf{0}^{(E\times B_{H})}$; $\mathbf{m}_{\text{max}}\leftarrow-\infty^{(B_{H})}$; $\mathbf{s}_{\text{exp\_sum}}\leftarrow\mathbf{0}^{(B_{H})}$; $c_{\text{scale}}\leftarrow s_{\text{total}}\cdot s_{Q}\cdot s_{K}$
13:
14$\triangleright$ Load query factors (fixed for this program as N=1)
15:Load $\mathbf{A}_{Q,\text{local}}^{(R_{Q}\times B_{H})}$ from $\mathbf{A}_{Q}[b,0,h_{\text{start}}\dots,:]$
16:Load $\mathbf{B}_{Q,\text{local}}^{(D\times R_{Q})}$ from $\mathbf{B}_{Q}[b,0,:,:]$triangleright Dimensions may be transposed after loading for matmul
17:
18$\triangleright$ Iterate over $M_{\text{chunk}}$-sized chunks within the K/V block
19:for $m_{\text{chunk\_start}}$ from $m_{\text{block}\_\text{start}}$ to $m_{\text{block}\_\text{end}}-1$ step $M_{\text{chunk}}$ do
20:    $m_{\text{chunk\_\text{end}}}\leftarrow\min(m_{\text{chunk\_start}}+M_{\text{chunk}},m_{\text{block}\_\text{end}})$
21:    $M_{\text{curr}\_\text{chunk}}\leftarrow m_{\text{chunk\_\text{end}}}-m_{\text{chunk\_start}}$
22:    $\triangleright$ Load K/V factors for the current chunk
23:    Load $\mathbf{a}^{K}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},B_{H})$; $\mathbf{a}^{V}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},B_{H})$; $\bm{b}^{K}_{\text{chunk}}(M_{\text{curr}\_\text{chunk}},D)$; $\bm{b}^{V}_{\text{chunk}}(E,M_{\text{curr}\_\text{chunk}})$triangleright Layouts optimized for memory access and matmuls
24:    $\bm{b}^{V}_{\text{chunk}}\leftarrow\bm{b}^{V}_{\text{chunk}}\cdot s_{V}$
25:    $\triangleright$ Core TPA Score Calculation for the chunk
26:    $S1_{\text{chunk}}\leftarrow\text{MatMul}(\bm{b}^{K}_{\text{chunk}},\mathbf{B}_{Q,\text{local}})$triangleright Shape: $(M_{\text{curr}\_\text{chunk}},R_{Q})$
27:    $S2_{\text{chunk}}\leftarrow\text{MatMul}(S1_{\text{chunk}},\mathbf{A}_{Q,\text{local}})$triangleright Shape: $(M_{\text{curr}\_\text{chunk}},B_{H})$
28:    $S3_{\text{chunk}}\leftarrow S2_{\text{chunk}}\odot\mathbf{a}^{K}_{\text{chunk}}\cdot c_{\text{scale}}$triangleright Shape: $(M_{\text{curr}\_\text{chunk}},B_{H})$
29:    $\triangleright$ Online Softmax Update for the chunk
30:    $\mathbf{m}_{\text{max\_local}}\leftarrow\max_{axis=0}(S3_{\text{chunk}})$triangleright Shape: $(B_{H})$
31:    $\mathbf{m}_{\text{max\_new}}\leftarrow\max(\mathbf{m}_{\text{max}},\mathbf{m}_{\text{max\_local}})$
32:    $\mathbf{p}_{\text{num}}\leftarrow\exp(S3_{\text{chunk}}-\mathbf{m}_{\text{max\_new}}[\text{None},:])$
33:    $\mathbf{s}_{\text{exp\_sum\_local}}\leftarrow\sum_{axis=0}(\mathbf{p}_{\text{num}})$
34:    $\mathbf{p}_{\text{weighted\_av}}\leftarrow(\mathbf{p}_{\text{num}}/\mathbf{s}_{\text{exp\_sum\_local}}[\text{None},:])\odot\mathbf{a}^{V}_{\text{chunk}}$
35:    $\mathbf{o}_{\text{chunk}}\leftarrow\text{MatMul}(\bm{b}^{V}_{\text{chunk}},\mathbf{p}_{\text{weighted\_av}})$triangleright Shape: $(E,B_{H})$
36:    $\triangleright$ Update global (M-block level) accumulators
37:    $\mathbf{s}_{\text{exp\_sum\_prev\_rescaled}}\leftarrow\mathbf{s}_{\text{exp\_sum}}\cdot\exp(\mathbf{m}_{\text{max}}-\mathbf{m}_{\text{max\_new}})$
38:    $\mathbf{s}_{\text{exp\_sum}}\leftarrow\mathbf{s}_{\text{exp\_sum\_prev\_rescaled}}+\mathbf{s}_{\text{exp\_sum\_local}}$
39:    $\text{ratio}\leftarrow\mathbf{s}_{\text{exp\_sum\_local}}/\mathbf{s}_{\text{exp\_sum}}$triangleright This is $\mathbf{s}_{\text{exp\_sum\_local}}/\mathbf{s}_{\text{exp\_sum\_new}}$
40:    $\mathbf{o}_{\text{accum}}\leftarrow(1-\text{ratio})\cdot\mathbf{o}_{\text{accum}}+\text{ratio}\cdot\mathbf{o}_{\text{chunk}}$
41:    $\mathbf{m}_{\text{max}}\leftarrow\mathbf{m}_{\text{max\_new}}$
42:end for
43:
44$\triangleright$ Store partial results for this program’s (batch, head_block, M_block)
45:Store $\mathbf{o}_{\text{accum}}$ into $\mathbf{O}_{\text{partial}}[b,p_{id_{M}},0,h_{\text{start}}\dots,:]$
46:$\mathbf{LSE}_{\text{val}}\leftarrow\log(\mathbf{s}_{\text{exp\_sum}})+\mathbf{m}_{\text{max}}$
47:Store $\mathbf{LSE}_{\text{val}}$ into $\mathbf{LSE}_{\text{partial}}[b,p_{id_{M}},h_{\text{start}}\dots]$
D.3 Additional Experimental Results
The following figures present additional speed comparisons for different embedding dimensions, with $d_{h}=64$ maintained. The y-axis represents $\log_{2}(\text{time})$ in seconds (lower is faster), and the x-axis represents $\log_{2}(\text{sequence length})$.
Detailed Analysis of Figure 5 (Main Text; Embedding Dimension 2048): Figure 5 in the main paper depicts speed comparisons for an embedding dimension of 2048. The results indicate that FlashTPA (blue line) is highly competitive. Across all tested batch sizes (1 to 16) for $d_{\text{model}}=2048$:
  • MHA (orange line) is consistently the slowest mechanism, with its decoding time increasing most rapidly with sequence length.
  • MQA (green line) and GQA (red line) offer significant speedups over MHA and perform very similarly to each other, often overlapping in the plots.
  • MLA (purple line) demonstrates strong performance, generally being faster than MQA/GQA, particularly at longer sequence lengths.
  • FlashTPA shows excellent scalability. While at very short sequence lengths (e.g., $2^{12}$ to $2^{13}$), its performance is comparable to MQA/GQA and MLA, its decoding time increases at a notably slower rate with sequence length. Consequently, FlashTPA becomes significantly faster than MQA/GQA for sequences longer than approximately $2^{14}$.
  • Compared to MLA, FlashTPA is consistently among the top two performers. In many instances, particularly at sequence lengths greater than $2^{14}$ or $2^{15}$, FlashTPA matches or slightly surpasses MLA in speed. The logarithmic scale for time suggests that these differences can be substantial in practice for very long contexts. For example, at a sequence length of $2^{19}$ across various batch sizes, FlashTPA often shows a visible advantage over MLA.
Figure 7 (Embedding Dimension 3072): With a larger embedding dimension of 3072, the relative performance trends observed in Figure 5 largely persist.
  • FlashTPA (blue line) remains one of the most efficient decoding methods. MHA (orange line) is consistently the slowest, while MQA (green line) and GQA (red line) offer considerable improvements over MHA.
  • MLA (purple line) and FlashTPA are the top two performers. FlashTPA consistently matches or exceeds the speed of MLA, particularly at longer sequence lengths (e.g., beyond $2^{15}$ or $2^{16}$ depending on the batch size). Its advantage often becomes more pronounced at the longest sequences tested ($2^{19}$). For instance, in batch size 1, TPA is clearly faster than MLA for sequence lengths $2^{16}$ and above. A similar trend is seen across other batch sizes, where TPA maintains a competitive edge or becomes superior at longer contexts.
This suggests that FlashTPA’s efficiency is well-maintained even as the model’s embedding dimension increases.
Figure 8 (Embedding Dimension 1024): For a smaller embedding dimension of 1024, similar trends are observed:
  • FlashTPA (blue line) is highly competitive. MHA (orange line) remains the least performant. MQA (green line) and GQA (red line) are faster than MHA.
  • At very short sequence lengths (around $2^{12}$ to $2^{13}$), MQA/GQA can be slightly faster than or comparable to TPA and MLA, especially for smaller batch sizes (e.g., Batch Size 1).
  • However, as sequence length increases, both MLA (purple line) and FlashTPA demonstrate superior scalability. FlashTPA generally matches or outperforms MLA, particularly for sequences longer than $2^{15}$. For example, with a batch size of 16, TPA shows a clear speed advantage over MLA for sequence lengths $2^{16}$ and greater.
These results across different embedding dimensions highlight the robustness of FlashTPA’s decoding speed advantages, especially for long sequences where it consistently ranks as one of the fastest, if not the fastest, attention mechanisms among those tested.
(a) Batch Size=1
(b) Batch Size=2
(c) Batch Size=4
(d) Batch Size=8
(e) Batch Size=16
Figure 7: Decoding time comparison of different attention mechanisms with an embedding dimension of 3072 and $d_{h}=64$.
(a) Batch Size=1
(b) Batch Size=2
(c) Batch Size=4
(d) Batch Size=8
(e) Batch Size=16
Figure 8: Decoding time comparison of different attention mechanisms with an embedding dimension of 1024 and $d_{h}=64$.
Appendix E Higher-Order Tensor Product Attention
All prior discussions have focused on TPA where the query, key, and value matrices (e.g., $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$) are formed as a sum of $R_{Q}$ components. Each component is an outer product of two context-dependent vectors, one spanning the head dimension ($\mathbb{R}^{h}$) and the other spanning the feature-per-head dimension ($\mathbb{R}^{d_{h}}$), as detailed in Section 3.1 (e.g., $\mathbf{Q}_{t}=\frac{1}{R_{Q}}\mathbf{A}_{Q}(\mathbf{x}_{t})^{\top}\mathbf{B}_{Q}(\mathbf{x}_{t})$ implies $\mathbf{Q}_{t}=\sum_{r}\mathbf{a}_{r}\bm{b}_{r}^{\top}$ where $\mathbf{a}_{r}$ are columns of $\mathbf{A}_{Q}^{\top}$ and $\bm{b}_{r}^{\top}$ are rows of $\mathbf{B}_{Q}$). We now generalize this by introducing additional latent factors in the construction of the feature-per-head vectors, leading to what we term higher-order TPA. This approach allows for more complex interactions in forming these feature vectors.
For instance, in a third-order factorization, the query tensor $\mathbf{Q}_{t}\in\mathbb{R}^{h\times d_{h}}$ for a single token $t$ is constructed as:
@@ Appendix E Higher-Order Tensor Product Attention
where $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{h}$. The term $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}}$ and the newly introduced factor $\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}}$ first form a matrix $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\otimes\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}\times d_{c}}$ via an outer product (as defined in Section 2). This matrix is then vectorized by $\operatorname{vec}(\cdot)$ into a column vector of dimension $d_{h}=d_{b}d_{c}$. The final query $\mathbf{Q}_{t}$ is formed by the sum of outer products between $\mathbf{a}^{Q}_{r}(\mathbf{x}_{t})$ and these resulting $d_{h}$-dimensional vectors. Analogous expansions apply to $\mathbf{K}_{t}$ and $\mathbf{V}_{t}$.
The additional factor $\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})$ can be viewed as a learnable, context-dependent modulation or gating term for the features generated by $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})$.
$\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{b}},\quad\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})\in\mathbb{R}^{d_{c}},\quad d_{h}=d_{b}d_{c}.$
This higher-order construction can enhance expressiveness. While introducing $\mathbf{c}^{Q}_{r}$ increases the parameter count for the factors, it might allow for the use of smaller base ranks ($R_{Q},R_{K},R_{V}$) to achieve comparable representational power, thus offering a different design choice. One could also explore tying or sharing $\mathbf{c}^{Q}_{r}$ across queries, keys, and values to manage parameter overhead.
From a memory perspective during inference, higher-order TPA maintains the benefit of factorized KV caching. Only the constituent factors $\mathbf{a}_{K}(\mathbf{x}_{t}),\mathbf{b}_{K}(\mathbf{x}_{t}),\mathbf{c}_{K}(\mathbf{x}_{t})$ (and similarly for values) for each past token need to be stored. A trade-off arises between model capacity and the overhead of memory and computation. Higher-order tensor decompositions can provide additional flexibility and potentially increased capacity.
E.1 RoPE Compatibility in Higher-Order TPA
Rotary positional embeddings (RoPE) remain compatible with higher-order factorizations. In second-order TPA, RoPE applies rotations to the $d_{h}$-dimensional feature vectors. This compatibility extends to higher-order TPA. Consider the case where RoPE is intended to primarily rotate feature pairs derived from the $\mathbf{b}^{Q}_{r}(\mathbf{x}_{t})$ components, while the structural influence of $\mathbf{c}^{Q}_{r}(\mathbf{x}_{t})$ components on the $d_{h}$-dimensional vector is preserved. More formally, RoPE acts on the $d_{h}$-dimensional vector $\operatorname{vec}(\mathbf{b}_{r}^{Q}\otimes\mathbf{c}_{r}^{Q})$ such that the transformation is equivalent to rotating $\mathbf{b}_{r}^{Q}$ to $\widetilde{\mathbf{b}}_{r}^{Q}=\mathbf{R}_{t}\mathbf{b}_{r}^{Q}$ (where $\mathbf{R}_{t}$ is the RoPE rotation matrix for $d_{b}$ dimensions) and then forming $\operatorname{vec}(\widetilde{\mathbf{b}}_{r}^{Q}\otimes\mathbf{c}_{r}^{Q})$. This is achieved by a specific RoPE transformation matrix $\mathbf{T}_{t}$ acting on the full $d_{h}$-dimensional vector, as stated in the following theorem.
Theorem E.1 (RoPE Compatibility in Higher-Order TPA).